# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ III-A1 Rule-Based Models
The intuitive clarity of rule-based models makes them highly interpretable and understandable. Their straightforward structure eliminates the need for post-hoc analysis, which is often required for more complex models. Furthermore, a rule wrapper in these models encapsulates key information, making it accessible to a non-expert audience and enabling it to operate effectively as a standalone prediction model. The transparency in rule-based models extends their utility beyond their standalone capabilities. They are also used to clarify the predictions of more intricate models by generating and applying rules to link sophisticated ML techniques with approachable interpretability.
@@ III-A2 Regression Models
Linear and logistic regression models stand as two important regression models. Linear Regression is often considered the foundational regression technique, distinguished by its straightforward nature and high linearity that predicts outcomes as a direct sum of input features [106]. The weight of the coefficient of linear regression is easy to quantify and interpret, which is why it is used in various fields to explain the predictions. Its simplicity in quantifying and interpreting the weight of coefficients makes Linear Regression a popular choice across various fields. These models are resistant to overfitting, though they face challenges in accurately representing feature relationships and are significantly influenced by outliers on regression boundaries and results.
On the other hand, logistic regression, a derivative of linear regression, serves as a classification model. It advances the basic regression approach by calculating probabilities for classification tasks and assigning weights to features that culminate in a definitive value range between 0 and 1.
This model’s strength lies in its ability to provide probabilities alongside classifications which offers a nuanced view of outcomes. Lundberg’s research [73] further enhances this model’s utility by integrating logistic regression with gradient-boosted trees for predicting synthetic labels and augmenting the explainability of tree-based models. Despite their transparency, these regression models often require additional post-hoc explainability tools, like visual aids, to make their predictions accessible to those not well-versed in statistical methodologies.
@@ III-A3 Decision Trees
Decision trees are structured hierarchically, segmenting data into various subsets based on distinct features. The terminal subsets, known as leaf nodes, play a crucial role in resolving classification and regression challenges. These trees offer transparent models that enable domain experts to understand how they work. Furthermore, the exploration of these trees can lead to the discovery of new relationships and insights. Blanco-Justicia and Domingo-Ferrer [18] leverage decision trees as surrogate models to elucidate black-box models, constructing these trees from segmented portions of the training dataset. This approach assumes that the person responsible for providing explanations has access to the training data and the black-box model. Nevertheless, decision trees encounter scalability issues with large datasets in real-world applications, which diminish their explainability as the tree complexity increases. This complexity requires the adoption of post-hoc explainability methods to maintain clarity.
@@ III-A4 Bayesian Models
Bayesian models excel in providing a high degree of interpretability and explainability, offering nuanced insights into the statistical interplay between variables. This capability makes them particularly adept for applications where clear, comprehensible explanations are essential, such as demonstrating the correlation between diseases and their symptoms. Hence, In the realm of medical research, the application of Bayesian methods has been notably effective. For instance, Arrieta et al. [11] demonstrate the utility of Bayesian approaches in healthcare analytics, highlighting their potential to the complex relationships within medical data. Similarly, the Naive Bayes classifier, as discussed by Rana et al. [93], serves as a robust algorithm for predictive modeling. This classifier efficiently tackles both binary and multi-classification problems by calculating the probabilities of individual elements, subsequently employing Bayes’ theorem to identify the most probable outcome. Moreover, to enhance explainability, Ren et al. [94] introduce an approach that integrates a neural network with an explainable Bayesian network. This novel method yields more accurate predictions and provides intuitive explanations for diagnoses.
@@ III-A5 K-Nearest Neighbour
The KNN algorithm is a paradigm of transparent, interpretable models utilized for both classification and regression tasks. At its core, KNN operates on a simple yet effective principle: for classification, it determines a test sample’s class based on the majority vote from its nearest neighbors, and for regression, it computes the average outcome of these neighbors. The selection of an appropriate K—the number of neighbors to consider— emerges as a critical factor, which can influence the model’s ability to accurately assess the proximity between data points and thus define the neighborhood essential for prediction.
The interpretability of KNN is significantly influenced by the chosen features, the distance metric utilized, and the number of neighbors. While models with extensive features may obscure interpretability, a KNN model characterized by a concise, well-selected feature set remains one of the main models for interpretable outcomes. This duality highlights KNN’s utility, especially in applications where the model’s transparency is important. In a study, Aslam et al. [13] showcase the application of various supervised ML-based models, including KNN, utilizing XAI techniques. These models provide crucial, timely medical insights within a fetal health monitoring system, illustrating KNN’s adaptability and its capacity for clear, interpretable insights, even in complex domains.
@@ III-A6 Generalized Linear and Additive Models
The Generalized Additive Model (GAM) represents a notable advancement in statistical modeling, combining the benefits of linearity and interpretability. This model assigns values to variables by integrating numerous undefined functions specific to regression models and enhancing accuracy without compromising on interpretability. One of the unique features of GAM is its ability to allow users to evaluate the significance of each variable by examining its impact on the predicted outcome. Notably, GAM models exhibit algorithmic transparency and are regarded as simulatable due to their minimal dimensionality issues. Furthermore, the Generalized Linear Model (GLM) broadens the scope of linear regression by being interpretable, preserving the feature weights combination, and capturing non-Gaussian relationships and dependencies.
@@
Despite the inherent transparency and explainability of such models, there is an ongoing exploration of undirected graphical models to enhance the model’s trustworthiness further. It is recognized that transparency alone may not suffice to ensure straightforward explainability. As models evolve, certain complexities may emerge, rendering them less interpretable. This complexity necessitates the development of post-hoc explanations to maintain the clarity and reliability of the models.
@@ III-B Opaque ML-based Models
III-B1 Random Forest (RF)
RFs are composed of numerous individual decision trees, each of which contributes to the model’s predictions by dividing the input space into smaller segments and averaging the outcomes across these segments. As the complexity of the problem increases, the number of decision trees within the forest grows exponentially, enhancing the model’s ability to address complex issues but simultaneously diminishing its explainability. RFs were developed to mitigate the overfitting issues associated with single decision trees and to improve their accuracy, with the overall prediction being the mean of the outputs from numerous trees. This strategy of combining multiple trees aims to reduce the variance of the final model. Each tree is trained on a unique subset of the dataset, allowing for an aggregated prediction that incorporates varied insights. However, the complexity inherent in the RF model complicates direct interpretation, leading to the necessity for post-hoc explainability techniques to unravel the decision-making process. In their study, Zhao et al. [129] propose a visual analytic system that explores various interpretive perspectives for analyzing and explaining the predictions generated by RFs, offering a comprehensive approach to understanding these models.
III-B2 Support Vector Machine (SVM)
A Support Vector Machine (SVM) is fundamentally based on geometrical principles, initially conceived for linear classification and subsequently extended to non-linear scenarios. In essence, an SVM constructs a hyperplane or a set of hyperplanes within a high or infinite-dimensional space, serving purposes across classification, regression, outlier detection, and even clustering tasks. A hyperplane achieves optimal separation when it maximizes the distance to the nearest point of the training dataset, as a larger margin correlates with a lower generalization error of the classifier. Owing to their remarkable predictive and generalization capabilities, SVMs are among the most prevalently utilized ML models. However, due to their complex dimensionality, they are often regarded as opaque, making their decision-making process less transparent. In their research, Vieira and Digiampietri [116] explore the use of decision trees to derive rules from SVMs, which provides explanations for the classifications made by SVM classifiers and enhances their interpretability.
@@ III-B3 Multi-Layer Neural Network
Numerous ML methodologies, including neural networks, have been specifically designed for DL applications. These models are computationally intensive but provide unparalleled performance across a wide range of applications. Neural networks are inherently considered black-box models due to their complex internal mechanisms, which require post-hoc explainability efforts to elucidate their decision-making processes. In the work referenced by Sharma et al. [104], the focus is on utilizing a multi-layer perceptron neural network for the risk prediction of default loans, with the explanation of model decisions facilitated through a sensitivity analysis technique.
A wide range of explainability techniques has been developed specifically for neural networks to enhance the transparency of their operations. In the following sections, we will delve into these techniques, providing a clear understanding of the methods used to interpret the complex functions of neural networks.
@@ III-C Model-Agnostic Techniques for Post-Hoc Explainability
In the upcoming section, we will explore various methods that necessitate post-hoc explainability. This need arises when an ML model lacks transparency and its decision-making process is not directly interpretable. To show the decision processes of these models, a separate post-hoc technique is applied. In other words, these post-hoc methods are designed to provide users with clear, understandable information about the predictions made by the model. Hence, they are particularly valuable for interpreting models that are considered black box models. The domain of model agnostic interpretability is divided into three main categories: Global Explanation, Local Explanation, and Visual Explanation. Global explanations aim to elucidate the model’s overall predictive behavior, Local explanations focus on clarifying individual predictions, and visual explanations seek to create easily understandable visual representations of the model’s functions.
@@ III-C1 Global Explanation
It refers to the process of achieving explainability at a comprehensive level, where the focus is on understanding the features that significantly influence the predictions of specific models. In scenarios where the prediction process is not inherently interpretable, a global surrogate model is employed. This model aims to replicate the predictions of the complex black-box model using a more interpretable ML-based approach. It is often termed the approximate or meta-model; the global surrogate model is a simpler stand-in for the original complex model. By analyzing this surrogate model, insights can be drawn about the underlying mechanisms of the black-box model. For instance, Islam et al. [51] demonstrated this concept by using the CART (Classification and Regression Trees) to approximate a random forest model’s behavior.
The surrogate model’s versatility is a key advantage. It allows for the evaluation of performance that closely aligns with the original, complex model. If the surrogate achieves similar accuracy, it can potentially eliminate the necessity for the original black-box model. Moreover, it is possible to develop several surrogate models for a single black-box model, each providing unique insights and interpretations. This method is efficient in its implementation and facilitates the explanation of sophisticated model dynamics in a clear and understandable manner.
@@ III-C2 Local Explanation
The scope of interpretability includes explaining the reasons for individual predictions of why a model made a specific decision in a given instance. This section will discuss local interpretability methods such as LIME (Local Interpretable Model-agnostic Explanations), KernalSHAP (SHapley Additive exPlanations), Shapley values, counterfactual explanations, and Logic Explained Networks (LENs).
LIME. For the first time, Ribeiro et al. [95] introduced a novel local explainability technique known as LIME. This method operates as a local surrogate model, generating interpretable predictions by approximating how the model behaves in the vicinity of a given prediction. It is designed to be model-agnostic, which makes it versatile across different ML models. Specifically, LIME focuses on elucidating the rationale behind individual predictions made by opaque, black-box models.
@@
In a practical application of LIME, Magesh et al. [75] utilizes this technique to interpret the predictions of a CNN model designed for the early detection of Parkinson’s disease. This study demonstrates the potential of LIME to provide valuable insights into real-world scenarios.
KernalSHAP. Among the various local interpretability methods developed, a significant challenge lies in determining the most suitable method for specific scenarios. To address this, Lundberg and Lee [74] proposed Shapley Additive exPlanations (SHAP), a concept derived from game theory that evaluates the importance of each feature in contributing to a particular prediction. The SHAP framework establishes a new class of additive feature importance measures characterized by a unique solution that exhibits desirable attributes.
@@
Shapley Values. Shapley values originate from coalitional game theory, framing each feature value of an instance as a ”player” and the prediction outcome as the ”payout.” This approach assigns a quantifiable contribution to each feature, which can demonstrate how significantly each one influences the final prediction. Shapley values are distinguished by key principles such as consistency and local accuracy. These principles ensure that the allocation of importance to features is both fair and interpretable, accurately reflecting each feature’s contribution to the outcome.
Counterfactual Explanations. Counterfactual explanations provide a compelling approach for local interpretation. This method stands out for its simplicity in implementation, as it does not need access to the underlying data or model. Counterfactual explanations focus on identifying which features would need alteration to achieve a specific desired outcome, thereby elucidating the reasoning behind model predictions. These explanations are particularly user-friendly because they illustrate how minimal changes in features can influence predictions. Nonetheless, one limitation of this method is its difficulty in accommodating categorical data across different levels. Related to the Counterfactual explanation, Molnar [85] discusses their application in models generating continuous predictions that showcase their utility in providing clear and actionable insights.
LENs LENs enhance the interpretability of neural networks by utilizing human-understandable predicates as inputs and translating predictions into First-Order Logic (FOL) explanations. These networks are highly adaptable and can function effectively in both supervised and unsupervised learning contexts. LENs can serve as direct classifiers, providing explanations for their predictions, or they can work alongside black-box classifiers to make their decisions interpretable. The learning process for LENs involves associating specific input features with output classes in supervised scenarios and generating logic rules that explain the conditions for predictions. In unsupervised learning, LENs identify patterns and relationships within the data, clustering similar data points and generating explanations that describe these clusters. Additionally, LENs can mimic the outputs of black-box models while generating FOL explanations, which leads to elucidating the decision-making process for these complex models [28].
@@ III-C3 Visual Explanation
Kramer et al. [58] demonstrated the application of ALE plots within the realm of real estate, employing them to discern which features significantly impact property values. This use case underscores the utility of ALE plots in practical, real-world analysis. Additionally, many researchers have adopted ALE plots as a method for visually exploring the nature of relationships between variables, assessing whether these relationships are linear or exhibit more complexity. This visual representation technique thus serves as a powerful tool for understanding and interpreting the effects of features on model predictions.

### v2

@@ III-A1 Rule-Based Models
The clarity of rule-based models makes them highly interpretable and understandable. Their straightforward structure eliminates the need for post-hoc analysis. They are also used to clarify the predictions of more intricate models by generating and applying rules to link sophisticated ML-based techniques with approachable interpretability.
@@ III-A2 Regression Models
Linear and logistic regression models stand as two important regression models. The weight of the coefficient of linear regression is easy to quantify and interpret, which is why it is used in various fields to explain the predictions. On the other hand, logistic regression strength lies in its ability to provide probabilities alongside classifications, which offers a nuanced view of outcomes.
Lundberg’s research [48] further enhances this model’s utility by integrating logistic regression with gradient-boosted trees for predicting synthetic labels and augmenting the explainability of tree-based models. Despite their transparency, these regression models often require additional post-hoc explainability tools, like visual aids, to make their predictions accessible to those not well-versed in statistical methodologies.
@@ III-A3 Decision Trees
Decision trees offer transparent models that enable domain experts to understand how they work. Furthermore, the exploration of these trees can lead to the discovery of new relationships and insights. Blanco-Justicia and Domingo-Ferrer [44] leverage decision trees as surrogate models to elucidate black-box models, constructing these trees from segmented portions of the training dataset.
This approach assumes that the person responsible for providing explanations has access to the training data and the black-box model. Nevertheless, decision trees encounter scalability issues with large datasets in real-world applications, which diminish their explainability as the tree complexity increases. This complexity requires the adoption of post-hoc explainability methods to maintain clarity.
@@ III-A4 Bayesian Models
Bayesian models excel in providing a high degree of interpretability and explainability, offering insights into the statistical interplay between variables. This capability makes them particularly useful for applications where clear, comprehensible explanations are essential, such as demonstrating the correlation between diseases and their symptoms.
Hence, in the realm of medical research, the application of Bayesian methods has been notably effective. For instance, Arrieta et al. [41] demonstrate the utility of Bayesian approaches in healthcare analytics, highlighting their potential to the complex relationships within medical data. Similarly, the Naive Bayes classifier, as discussed by Rana et al. [49], serves as a robust algorithm for predictive modeling. This classifier efficiently tackles both binary and multi-classification problems by calculating the probabilities of individual elements, subsequently employing Bayes’ theorem to identify the most probable outcome.
@@
The k-NN algorithm operates on a simple yet effective principle: for classification, it determines a test sample’s class based on the majority vote from its nearest neighbors, and for regression, it computes the average outcome of these neighbors. The interpretability of k-NN is significantly influenced by the chosen features, the distance metric utilized, and the number of neighbors. While models with extensive features may obscure interpretability, a k-NN model characterized by a concise, well-selected feature set remains one of the main models for interpretable results. In a study, Aslam et al. [50] showcase the application of various supervised ML-based models, including k-NN, utilizing XAI techniques.
@@ III-A6 Generalized Linear and Additive Models
The Generalized Additive Model (GAM) represents an advancement in statistical modeling, combining the benefits of linearity and interpretability. This model assigns values to variables by integrating numerous undefined functions specific to regression models and enhancing accuracy without compromising interpretability. One of the unique features of GAM is its ability to allow users to evaluate the significance of each variable by examining its impact on the predicted outcome. GAM models exhibit algorithmic transparency and are regarded as simulatable due to their minimal dimensionality issues.
@@
Despite the inherent transparency and explainability of these models, ongoing research explores undirected graphical models to further enhance their trustworthiness. Transparency alone may not always guarantee straightforward explainability, as increasing model complexity can reduce interpretability. This necessitates the development of post-hoc explanations to maintain clarity and reliability. As mentioned above and illustrated in Figure 3, the dashed arrow represents cases where even transparent models may require post-hoc explanations in complex scenarios. This highlights the dynamic nature of explainability, where certain conditions still necessitate additional interpretability techniques to ensure model reliability and trust.
@@ III-B Opaque ML-based Models
III-B1 Random Forest
Random Forests (RFs) consist of multiple decision trees, each dividing the input space into smaller segments and averaging outcomes. As problem complexity increases, more trees are needed, improving accuracy but reducing explainability. RFs were designed to mitigate overfitting in single decision trees by averaging predictions across multiple trees, reducing variance. Each tree is trained on a unique data subset, ensuring diverse insights.
However, the model’s complexity necessitates post-hoc explainability techniques. Zhao et al. [52] introduce a visual analytic system to enhance interpretability, offering a comprehensive approach to understanding RF predictions.
III-B2 Support Vector Machine
SVM constructs a hyperplane or a set of hyperplanes within a high or infinite-dimensional space, serving purposes across classification, regression, outlier detection, and even clustering tasks. A hyperplane achieves optimal separation when it maximizes the distance to the nearest point of the training dataset, as a larger margin correlates with a lower generalization error of the classifier. Owing to their remarkable predictive and generalization capabilities, SVMs are among the most widely utilized ML models. However, due to their complex dimensionality, they are often regarded as opaque, making their decision-making process less transparent.
Based on this, Vieira and Digiampietri [53] explore the use of decision trees to derive rules from SVMs, which provides explanations for the classifications made by SVM classifiers and enhances their interpretability.
@@ III-B3 Multi-Layer Neural Network
These models are computationally intensive but provide unparalleled performance across a wide range of applications. Neural networks are inherently considered black-box models due to their complex internal mechanisms. In the study by Sharma et al. [54], the focus is on utilizing a multi-layer perceptron neural network for the risk prediction of default loans, with the explanation of model decisions facilitated through a sensitivity analysis technique.
@@ III-C Model-Agnostic Techniques for Post-Hoc Explainability
Post-hoc explainability methods play a crucial role in interpreting complex ML-based models, especially in high-stakes applications where decision transparency is required. One way to classify post-hoc explainability methods is based on their dependency on the model structure. In this classification, we identify two groups: model-agnostic methods, which can be applied to any model, and model-specific methods, which are tailored to particular model architectures.
Model-agnostic techniques can be applied across various ML architectures, offering greater flexibility. However, this flexibility often comes at the cost of precision in explanations, as these methods approximate model behavior rather than providing direct interpretability. Given the increasing reliance on ML in security-sensitive domains such as malware detection, adopting robust and interpretable post-hoc techniques is essential to ensure trust and accountability. The domain of model agnostic interpretability is divided into three main categories: Global Explanation, Local Explanation, and Visual Explanation.
@@ III-C1 Global Explanation
Global explainability provides insights into how a model makes decisions across all instances, rather than focusing on individual predictions. This is essential for understanding which features are most influential and how they interact at a systemic level.
A widely used approach to achieving global explainability is the use of surrogate models, which approximate the decision-making process of a complex black-box model by training a more interpretable alternative (e.g., decision trees, linear models) on the same dataset. These models enable researchers to analyze feature contributions in a transparent manner, facilitating model interpretation.
However, surrogate models introduce trade-offs. Since they approximate rather than replicate the black-box model’s decision boundaries, they may introduce inaccuracies, particularly in the presence of nonlinear relationships or complex feature interactions. Despite these limitations, they remain a practical tool for obtaining high-level insights into opaque models.
In cybersecurity and malware detection, global explanations help identify critical risk factors in security assessments. For instance, a surrogate model trained on malware classification outputs can reveal whether API call sequences, file metadata, or network behaviors are the most significant predictors. This understanding allows security analysts to refine detection rules, improve feature selection, and enhance model robustness.
Surrogate models serve as interpretable stand-ins for complex models, offering insights into their underlying mechanisms. For example, Islam et al. [5] demonstrated the effectiveness of this approach by using Classification and Regression Trees (CART) to approximate a random forest’s decision-making process. If a surrogate achieves comparable performance, it may reduce reliance on the original model, particularly when interpretability is prioritized. Additionally, multiple surrogate models can be developed for a single black-box system, each providing distinct perspectives on model behavior. This approach enhances transparency and facilitates the comprehension of sophisticated decision processes.
@@ III-C2 Local Explanation
Local explainability focuses on understanding why a model makes a specific prediction for a single instance rather than explaining overall model behavior. These methods help users answer questions such as, ”Why was this particular file classified as malware?” or ”What features contributed to this anomaly?” Unlike global explanations, which provide an overview of feature importance across an entire dataset, local explanations offer insights into decision-making at an individual level.
Local explanation techniques are particularly valuable in high-stakes applications like cybersecurity, where understanding why a model flagged a file as malicious can assist analysts in investigating threats, identifying adversarial attacks, or refining detection rules. Additionally, local explanations play a critical role in bias detection and fairness assessments, ensuring that models do not make decisions based on unintended or discriminatory features.
Several widely used model-agnostic techniques exist for local explainability, including LIME (Local Interpretable Model-Agnostic Explanations), KernelSHAP (Shapley Additive Explanations), Shapley values, counterfactual explanations, and Logic Explained Networks (LENs). The following sections discuss these methods in more detail, highlighting their strengths, limitations, and practical applications.
LIME. For the first time, Ribeiro et al. [55] introduce a novel local explainability technique known as LIME. This method operates as a local surrogate model, generating interpretable predictions by approximating how the model behaves in the vicinity of a given prediction. It is designed to be model-agnostic, which makes it versatile across different ML-based models.
@@
In a practical application of LIME, Magesh et al. [56] utilize this technique to interpret the predictions of a CNN model designed for the early detection of Parkinson’s disease. This study demonstrates the potential of LIME to provide valuable insights into real-world scenarios.
KernalSHAP. Among the various local interpretability methods developed, a significant challenge lies in determining the most suitable method for specific scenarios. To address this, Lundberg and Lee [57] propose Shapley Additive exPlanations (SHAP), a concept derived from game theory that evaluates the importance of each feature in contributing to a particular prediction. The SHAP framework establishes a new class of additive feature importance measures characterized by a unique solution that exhibits desirable attributes.
@@
Shapley Values. They originate from coalitional game theory, framing each feature value of an instance as a ”player” and the prediction outcome as the ”payout.” This approach assigns a quantifiable contribution to each feature, which can demonstrate how significantly each one influences the final prediction. Shapley values are distinguished by key principles such as consistency and local accuracy. These principles ensure that the allocation of importance to features is both fair and interpretable, accurately reflecting each feature’s contribution to the outcome.
Counterfactual Explanations. Counterfactual explanations provide a compelling approach for local interpretation. This method stands out for its simplicity in implementation, as it does not need access to the underlying data or model. Counterfactual explanations focus on identifying which features would need alteration to achieve a specific desired outcome, thereby elucidating the reasoning behind model predictions. These explanations are particularly user-friendly because they illustrate how minimal changes in features can influence predictions. Nonetheless, one limitation of this method is its difficulty in accommodating categorical data across different levels. Related to the Counterfactual explanation, Molnar [58] discusses their application in models generating continuous predictions that showcase their utility in providing clear and actionable insights. In malware classification, counterfactual explanations can highlight the minimal changes required to alter a model’s decision, offering valuable insights for both threat analysis and adversarial defense strategies. This approach is particularly useful in detecting adversarial attacks, as it helps identify which modifications in malware features could evade detection, thereby strengthening model robustness.
LENs. They enhance the interpretability of neural networks by utilizing human-understandable predicates as inputs and translating predictions into First-Order Logic (FOL) explanations. These networks are highly adaptable and can function effectively in both supervised and unsupervised learning contexts. LENs can serve as direct classifiers, providing explanations for their predictions, or they can work alongside black-box classifiers to make their decisions interpretable. The learning process for LENs involves associating specific input features with output classes in supervised scenarios and generating logic rules that explain the conditions for predictions. In unsupervised learning, LENs identify patterns and relationships within the data, clustering similar data points and generating explanations that describe these clusters. Additionally, LENs can mimic the outputs of black-box models while generating FOL explanations, which leads to elucidating the decision-making process for these complex models [59]. LENs bridge the gap between symbolic reasoning and neural networks by incorporating logical constraints into the learning process, making their decisions more interpretable. Unlike traditional neural networks, which act as black boxes, LENs provide structured, rule-based explanations that enhance transparency and trust in AI-driven decision-making.
@@ III-C3 Visual Explanation
Kramer et al. [61] demonstrated the application of ALE plots within the realm of real estate, employing them to discern which features significantly impact property values. This use case underscores the utility of ALE plots in practical, real-world analysis. Additionally, many researchers have adopted ALE plots as a method for visually exploring the nature of relationships between variables, assessing whether these relationships are linear or exhibit more complexity.